OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=1, b=7, c=?

OpenStudy (anonymous):

are you familiar with this ? a^2+b^2=c^2

myininaya (myininaya):

\[a^2+b^2=c^2 =>c=\sqrt{a^2+b^2}\]

myininaya (myininaya):

plug in

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

oohh waiiit yeea ythagorean theorm

OpenStudy (saifoo.khan):

so did u got the asnwer?

OpenStudy (anonymous):

is it 50?

OpenStudy (saifoo.khan):

\[c^2 = 50\]

OpenStudy (saifoo.khan):

\[c = \sqrt{50}\]

OpenStudy (saifoo.khan):

\[c = 5 \sqrt{2}\]

OpenStudy (anonymous):

yaaaaayi almost ot itt

OpenStudy (saifoo.khan):

Rightt.

OpenStudy (anonymous):

kiss3fist - great job !